what value of y makes the two expressions equal 5y-8 3y+5
Answer by RAY100(1637)

Set (5y-8) =(3y+5)
subst (3y) each side,,and add (8) each side
2y=13
divide by 2 each side
y= 13/2 =6.5
,
check
5(6.5)-8 = 3(6.5)+5
32.5-8 = 19.5+5
24.4 = 24.5